The gravitational force between two sumo wrestlers just before they collide (r=0.0250m) is 2.770*10^-3 N.
If one wrestler has a mass of 157.0 kg, what is his opponents mass?
1
Expert's answer
2011-06-07T06:34:51-0400
As F = (G m1 m2) / r2 ,
m2 = (r2 *F) / (G m1) = (0.0252*2.77*10-3)/(6.673×10-11 *157) = 165.2 kg.